# Topological columnar nano-SQUID based on a 3D topological insulator

### v2

@@ II Results
III Discussion
We now show theoretically that the nano-SQUID formed by the TI top and bottom surface states placed between the superconducting leads realizes a topological superconducting state over an extended flux range from $\Phi=(n-\frac{1}{2})\Phi_{0}^{s}$ to $(n+\frac{1}{2})\Phi_{0}^{s}$ around odd $n$ of flux quanta piercing the nanowire, see Fig. 4b for a sketch.
Recall that the spin-momentum locking in the TI surface states prohibits normal reflection at the NS interface for the perpendicular mode ($k$=0) and guarantees a pair of perfectly transmitted, non-degenerate ABSs (even- and odd-parity ABS) in the short-junction limit [4]. They present $4\pi$ periodicity and cross zero-energy at $\varphi=(2m+1)\pi$ with integer $m$, where the fermion parity of the ground state alternates [46, 47]. When this $(2m+1)\pi$ crossing occurs in only one of the junctions in a columnar nano-SQUID, it guarantees the appearance of a 1D topological phase [48, 4]. Now let us first consider the special case when an odd number of flux quanta pierces the nanowire, which leads to the superconducting phase to wind by an odd multiple of $2\pi$ in the nano-SQUID. The $2\pi$ relative phase difference (for example, phase zero at the top and $2\pi$ at the bottom) implies that the fermion parity on the top and bottom surfaces is opposite, making the SQUID topological. In this situation, the two ends of the columnar TI nano-SQUID become a boundary between the topological and trivial regions, which implies that MZMs should appear at the ends (Fig. 4b).
@@ III Discussion
Our experiment demonstrates that the conceived columnar TI nano-SQUID can actually be fabricated with a new side-contacting technique and our device shows that the supercurrent is surface-dominated, which guarantees the prominent role of spin-momentum locking that dictates the appearance of perfectly-transmitted ABSs. Our theory shows that the occurrence of the topological phase transition in this nano-SQUID as a function of flux $\Phi$ is guaranteed by symmetry. Since the 1D topological superconductivity realized in a TI junction by phase-biasing is independent of chemical potential and the whole TI surface is gapped, our columnar nano-SQUID is a promising platform to host robust MZMs. To confirm the non-Abelian statistics of MZMs in this platform, it is important to realize a hard gap to completely get rid of low-energy quasiparticles and to theoretically develop concrete strategies to perform braiding.
@@ Methods
Theory: To treat the TINW, we use a model for the 3D TI material BSTS [2, 23, 24, 50] with finite square cross-section of side length $12L+$ 1, such that the Bogoliubov-de Gennes Hamiltonian is given by
@@
For simplicity we choose parameters $A=1$, $M=-1.5$, $B=0.5$, $L=4$, and $\mu=0.5$. We choose the pairing potentials $\Delta=0.02$, which is smaller than the subband gap, and $\Delta=0.25$, which is larger than the subband gap (see Fig. 4). We set $N$=3 such that only the side sites in the cross-section have an induced pairing potential. It should also be noted that, due to the finite extent of the wavefunctions into the nanowire, the value of flux corresponding to the normal-state flux quantum $\Phi=\Phi_{0}^{n}=h/e$ is obtained from the normal-state band structure (see supplement).
@@ Acknowledgments
This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant Agreement No. 741121) and was also funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - Cluster of Excellence Matter and Light for Quantum Computing (ML4Q) EXC 2004/1 - 390534769, as well as by the DFG under CRC 1238 - 277146847 (Subprojects A04 and B01). E.N. acknowledges support by the Studienstiftung des deutschen Volkes. The work at Massachusetts Institute of Technology was supported by a Simons Investigator Award from the Simons Foundation.

### v3

@@ II Columnar nano-SQUID
III Topological Superconductivity
We now show theoretically that the columnar TI nano-SQUID realizes a topological superconducting state over an extended flux range from $\Phi=(n-\frac{1}{2})\Phi_{0}^{s}$ to $(n+\frac{1}{2})\Phi_{0}^{s}$ around odd $n$ of flux quanta piercing the nanowire (see Fig. 4b for a sketch) and the only condition for the topological phase is a top/bottom asymmetry. Also, this topological phase is essentially insensitive to disorder. This prediction is the most important result of this paper.
Recall that the spin-momentum locking in the TI surface states prohibits normal reflection at the NS interface for the perpendicular mode ($k_{x}$ = 0) and guarantees a pair of perfectly transmitted, non-degenerate ABSs (even- and odd-parity ABS) in the short-junction limit [4]. They present $4\pi$ periodicity and cross zero-energy at $\varphi=(2m+1)\pi$ with integer $m$, where the fermion parity of the ground state alternates [44, 45]. When this $(2m+1)\pi$ crossing occurs in only one of the junctions in a columnar nano-SQUID, it guarantees the appearance of a 1D topological phase [46, 4]. Now let us first consider the special case when an odd number of flux quanta pierces the nanowire, which leads to the superconducting phase to wind by an odd multiple of $2\pi$ in the nano-SQUID. The $2\pi$ relative phase difference (for example, phase zero at the top and $2\pi$ at the bottom) implies that the fermion parity on the top and bottom surfaces is opposite, making the SQUID topological. In this situation, the two ends of the columnar TI nano-SQUID become a boundary between the topological and trivial regions, which implies that MZMs should appear at the ends (Fig. 4b).
@@ III Topological Superconductivity
IV Robustness against disorder
In experimental devices it is inevitable that disorder, e.g., in chemical potential and flux, will be present [42]. We analyzed the stability of the topological phase by numerically simulating the energy gap to excited states (see supplement for details). We find that a gapped topological phase exists over a large region of the parameter space. In particular, the phase shows considerable stability against variations in chemical potential, persisting for the full range of the TI bulk band gap. The phase is also maintained across a sizable range of both the phase difference and the magnetic flux. This large region of phase space where topological superconductivity should be realised strongly suggests that the topological phase has a significant tolerance to the sources of disorder that can be expected in our devices. The largest energy excitation gaps are achieved for chemical potentials close to the charge neutrality point, when a flux quantum threads the nanowire, and for a phase difference of $\pi$ across the junction. The fact that this regime can actually be achieved in the current devices corroborates the promise of this platform to host robust topological superconductivity.
V Conclusion and outlook
Our experiment demonstrates that the conceived columnar TI nano-SQUID can actually be fabricated with a new side-contacting technique and our device shows that the supercurrent is surface-dominated, which guarantees the prominent role of spin-momentum locking that dictates the appearance of perfectly-transmitted ABSs. Our theory shows that the occurrence of the topological phase transition in this nano-SQUID as a function of flux $\Phi$ is guaranteed by symmetry. Since the 1D topological superconductivity realized in a TI junction by phase-biasing is independent of chemical potential and the whole TI surface is gapped, our columnar nano-SQUID is a promising platform to host robust MZMs. An obvious next step is to experimentally confirm the emergence of the topological phase in the TI nano-SQUID, and a first experiment would be to detect a zero-bias conductance peak (ZBCP) due to MZMs with a tunnel probe. In this regard, it would be useful to theoretically investigate how a trivial ZBCP of non-MZM origin can appear in this platform. To get rid of trivial ZBCPs and to perform braiding to confirm the non-Abelian statistics of the MZMs, it will be important to realize a hard gap in this platform. Theoretical developments of concrete strategies to perform braiding in this promising platform would also be highly desirable.
@@ Methods
Theory: To treat the TINW, we use a model for the 3D TI material BSTS [2, 19, 20, 48] with finite square cross-section of side length $2L+1$ in the Kwant package [49], such that the Bogoliubov-de Gennes Hamiltonian is given by
@@
For simplicity we choose parameters $A=1$, M=1.5M=-1.5, $B=0.5$, $L=4$, and $\mu=0.5$. We choose the pairing potentials $\Delta=0.02$, which is smaller than the subband gap, and $\Delta=0.25$, which is larger than the subband gap (see Fig. 4). We set NN=3 such that only the side sites in the cross-section have an induced pairing potential. It should also be noted that, due to the finite extent of the wavefunctions into the nanowire, the value of flux corresponding to the normal-state flux quantum $\Phi=\Phi_{0}^{n}=h/e$ is obtained from the normal-state band structure (see supplement).
@@ Acknowledgments
We thank Oliver Breunig for the help in Fraunhofer-pattern fittings. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant Agreement No. 741121) and was also funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - Cluster of Excellence Matter and Light for Quantum Computing (ML4Q) EXC 2004/1 - 390534769, as well as by the DFG under CRC 1238 - 277146847 (Subprojects A04 and B01). E.N. acknowledges support by the Studienstiftung des deutschen Volkes. The work at Massachusetts Institute of Technology was supported by a Simons Investigator Award from the Simons Foundation.